OpenStudy (anonymous):

really need help with algebra 2

OpenStudy (anonymous):

What is the equation of a vertical line passing through the point (−4, 7)

OpenStudy (kamibug):

The equation of a vertical line always starts with "x=" because it only passes through one point on the x-axis. Just take the x-coordinate from the point you are given. :) The equation for this line would be x = -4.

OpenStudy (anonymous):

thank you so much. A few more questions please if you dont mind

OpenStudy (kamibug):



OpenStudy (kamibug):

Sure, I don't mind. :)

OpenStudy (anonymous):

On the graph of the equation 6x − y = −1, what is the value of the y-intercept?

OpenStudy (kamibug):

The y-intercept means where the line crosses the y-axis. Any point that touches the y-axis has a x-coordinate of 0. So, just plug in 0 for x in the equation and solve for y! :) Think you got this one from here?

OpenStudy (anonymous):

yintercept is -1?

OpenStudy (anonymous):

@KamiBug

OpenStudy (kamibug):

6(0) - y = -1 - y = -1 y = 1 <-- You have to get the opposite b/c we are looking for +y, not -y. :)

OpenStudy (anonymous):

ok thank you so much. a few more questions

OpenStudy (kamibug):

Okay:)

OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (−2, 6) and has a slope of −1?

OpenStudy (anonymous):

@KamiBug